Production increased by 145%. If the number of toys produced this year was 171,500, how many were produced last year?

Answer by Theo(13342)

171500 were produced this year.
that was a 145% increase over last year production.
if we let x equal last year production, then we get:
145% equals a multiplication factor of 1.45.
our equation for last year becomes:
x + 1.45x = 171500
if we combine like terms, we get:
2.45x = 171500
divide both sides of this equation by 2.45 to get:
x = 171500/2.45 which becomes:
x = 70,000
last year we produced 70,000
this year our production increased by 145%.
145% of 70,000 = 1.45 * 70,000 = 101,500
add 101,500 to 70,000 and you get 171,500 units produced this year.